a to the -5 times b to the 4 times c to the -2 times d. how do i rewrite this using positive eponents?
Answer by tutorcecilia(2152)

To convert negative exponents to positive exponents, take the reciprocal of the term with a positive exponent.
Example #1: a^-1 = a^-1/1 = 1/a^1
.
Example #2: b^-2 = b^-2/1 = 1/b^2
.
a^-5b^4c^-2d
.
=b^4d/a^5c^2
..
I took only the reciprocal of a^-5 and c^-2 because these were the only terms that contained a negative exponent.